When a potential difference V is applied between the ends of a wire of diameter d and length l , the current in the wire is I. What is the current when a potential difference of 2V is applied between the ends of a wire of the same material of diameter 2d and the length 2l ? Assume that the temperature of the wire remains constant. 


1
Expert's answer
2020-11-02T09:23:44-0500

resistance of wire:

"R = \rho\times\frac{l}{S}"

where "\rho" - electrical resistivity of material, l - length of wire, S - cross sectional area of wire.

current:

"I = \frac{V}{R}"

then:

"R = \frac{V}{I}"

this implies:

"\frac{V}{I} = \rho\times\frac{l}{S}"

Cross-sectional area for wire:

"S=\pi \times \frac{d^2}{4}"

this implies:

"\frac{V}{I} = \rho\times\frac{4\times l}{\pi\times d^2}"

expressing I:

"I=\frac{V\times \pi \times d^2}{4\times \rho \times l}"

expressing current for second case:

"I_{2}=\frac{2\times V\times \pi \times 4 \times d^2}{4\times \rho \times 2 \times l}"

simplifying

"I_{2}=\frac{V\times \pi \times d^2}{\rho \times l}"